# A 28nm Multiply-Accumulate ASIC Architecture for On-Chip Data Compression in MHz Frame Rate X-ray and Electron Pixel Detectors

### v1

@@ 3.1.4 Floating-Point Multipliers
Weights are typically represented using IEEE floating-point formats, such as IEEE FP32 or IEEE FP16. Floating-point multipliers are designed to perform the multiplication operations necessary for matrix multiplication between integer-valued pixel data and the floating-point weights. As illustrated in figure 4, the number of multipliers is equal to K × Number of Columns, while their size depends on both the pixel width and the weight width. The data from a single column is multiplied $K$ times with the corresponding weights stored in the SRAM. After this operation, all the results generated from the multiplier bank are sent to the corresponding accumulator.
@@ 5.2 Adjusting Weight Width
Note that while the weights now have narrower widths, we do not necessarily change the widths of the output of the multipliers. Since our input pixel values are unsigned 12-bit integers and our largest weight is of value 1, we want to be able to represent the same values as the pixel values in our output format. It turns out the products can all be represented in a standard 16-bit floating point format with only a slight loss in precision.
@@ 5.3 Adjusting Width of Accumulation Output
While adjusting the width of the weights has resulted in significant area savings, adjusting the accumulation width is motivated more by avoiding overflow. If we consider the the max pixel value of $2^{12}-1$, weights are between values -1 and 1, and that our design processes a 12×168 frame, the largest magnitude number that our accumulator can produces is $2^{12}\times 12\times 168=8255520$, which is greater than $2^{22}$. As a result, using IEEE half precision will result in an overflow, since it can only represent exponents between -15 and 15. If a value goes beyond these bounds, it will result in overflow. To account for this, we simply widen the output of the last level of adders in the accumulator to 17-bit values with e set to 6, extending the range of exponents to -31 to 31.
@@ 6 Logic Synthesis
Logic synthesis was performed on a $28\text{\,}\mathrm{nm}$ CMOS technology. Various partitioning schemes and compression ratios were implemented as part of a feasibility study for the current design. The synthesis process considered a pixel array of size 192 × 168, with pixel width of 12. Two design approaches were implemented. The first approach uses a constant data width of 16 bits for the weights, the output from the multiplier, and the final output of the accumulator. However, this method is not optimal in cases of overflow. The second approach adjusts the data width as needed, starting with 12-bit weights, a 16-bit output from the multiplier, and a 17-bit final accumulation output. This second method effectively saves area and prevents overflow; however, it sacrifices some precision.
@@ 8 Discussion
If we consider the case for K = 192, this corresponds to $74.3\text{\,}\mathrm{M}\mathrm{b}$ of weight data read from the SRAMs each microsecond. This results in an overall memory bandwidth of $74.3\text{\,}\mathrm{T}\mathrm{b}\mathrm{p}\mathrm{s}$, which is not feasible with current state-of-the-art off-chip memory technologies (e.g., HBM) due to the enormous number of interconnects required (over 100,000 interconnects in the case of K = 192 and FP12 floating-point weights). This substantial bandwidth significantly impacts the overall power consumption of the system.

### v2

@@ 3.1.4 Floating-Point Multipliers
Weights are typically represented using IEEE floating-point formats, such as IEEE FP32 or IEEE FP16. Floating-point multipliers are designed to perform the multiplication operations necessary for matrix multiplication between integer-valued pixel data and the floating-point weights. As illustrated in figure 4, the number of multipliers is equal to K × Number of Columns, while their size depends on both the pixel data width and the weight width. The data from a single column is multiplied $K$ times with the corresponding weights stored in the SRAM. After this operation, all the results generated from the multiplier bank are sent to the corresponding accumulator.
@@ 5.1 Floating Point Format
5.2 Floating Point Operator Implementation
The multipliers and adders used in the design are based on the IEEE 754 floating-point specification. However, we do not add support for infinity, NaN, or subnormals to reduce the complexity of the operators. The multipliers can be simplified due to one of the operands being an integer. As a result, the multiplier logic consists of an integer multiply, taking the mantissa of the weight and the full input value as operands. The leading one of the resulting product is determined through a tree of multiplexers, the output of which is fed into a shifter and rounded using round-to-even policy to produce the mantissa of the final output. The output of the multiplexer tree is also added to the exponent of the weight to determine the final output’s exponent value. Finally, since the input is an unsigned integer, the sign of the weight is passed through as the sign of the final output.
The adders use a comparator and multiplexers to determine which input has the largest magnitude. The mantissas are then normalized by rightward shifting the smaller magnitude number by the difference of the exponents. If the signs of the inputs match, then the normalized mantissas are added. Otherwise, the mantissa of the smaller magnitude number is subtracted from the other mantissa. In either case, the leading one of the sum is located to determine the new output exponent. The one’s position is also used to shift the sum before being rounded using round-to-even policy, resulting in the final output’s mantissa. The sign of the larger magnitude input is the sign of the final output.
@@ 5.3 Adjusting Weight Width
Note that while the weights now have narrower widths, we do not necessarily change the widths of the output of the multipliers. Since our input pixel values are unsigned 12-bit integers and our largest weight is of value 1, we want to be able to represent the same values as the pixel values in our output format. It turns out the products can all be represented in a standard 16-bit floating point format with only a slight loss in precision. The reasoning behind this is that the range of the floating point operand is [1,1][-1,1] while the range of the integer operand is $0,2^{12}-1]$, meaning the range of the product is $2^{12}+1,-2^{12}+1]$. FP16 can represent number of magnitude larger than $2^{12}$, justifying the need to set $e=5$ for the product floating point format. Furthermore, $m=10$ is chosen since an input value of high magnitude such as $2^{12}-1$ (i.e. 0b1111111111111111) will require an 11-bit mantissa to represent. While using 10-bits introduces loss, these extreme values are uncommon in our workloads thus the loss is negligible.
@@ 5.4 Adjusting Width of Accumulation Output
While adjusting the width of the weights has resulted in significant area savings, adjusting the accumulation width is motivated more by avoiding overflow. If we consider the max pixel value of $2^{12}-1$, weights are between values -1 and 1, and that our design processes a 12×168 frame, the largest magnitude number that our accumulator can produce is $2^{12}\times 12\times 168=8255520$, which is greater than $2^{22}$. As a result, using IEEE half precision will result in an overflow, since it can only represent exponents between -15 and 15. If a value goes beyond these bounds, it will result in overflow. To account for this, we simply widen the output of the last level of adders in the accumulators to 17-bit values with e set to 6, extending the range of exponents to -31 to 31. However, this overflow does not occur when calculating the partial sum using the adder tree, so the adder tree is entirely composed for 16-bit adders.
@@ 6 Logic Synthesis
Logic synthesis was performed on a $28\text{\,}\mathrm{nm}$ CMOS technology. Various partitioning schemes and compression ratios were implemented as part of a feasibility study for the current design. The synthesis process considered a pixel array of size 192 × 168, with pixel data width of 12. Two design approaches were implemented. The first approach uses a constant data width of 16 bits for the weights, the output from the multiplier, and the final output of the accumulator. However, this method is not optimal in cases of overflow. The second approach adjusts the data width as needed, starting with 12-bit weights, a 16-bit output from the multiplier, and a 17-bit final accumulation output. This second method effectively saves area and prevents overflow; however, it sacrifices some precision.
@@ 7 Physical Implementation
The design is scheduled to be submitted for tape-out, details of which will be presented in a forthcoming publication.
@@ 8 Discussion
If we consider the case for K = 192, this corresponds to 74.3 Mbits of weight data read from the SRAMs each microsecond. This results in an overall memory bandwidth of $74.3\text{\,}\mathrm{T}\mathrm{b}\mathrm{p}\mathrm{s}$, which is not feasible with current state-of-the-art off-chip memory technologies (e.g., HBM) due to the enormous number of interconnects required (over 100,000 interconnects in the case of K = 192 and FP12 floating-point weights). This substantial bandwidth significantly impacts the overall power consumption of the system.